OpenStudy (anonymous):

what is 6times 6 @saue7414

OpenStudy (midhun.madhu1987):

times means multiplied 6 times 6 = 6*6 =

OpenStudy (anonymous):

6*6=36

OpenStudy (anonymous):

thanks u

OpenStudy (anonymous):

You're welcome! @Ese_Muerte

OpenStudy (anonymous):

36 lOl